OpenStudy (yanasidlinskiy):

I just gotta question. In the quadratic Formula, like the ax^2+bx+c=0......Does it always have to =0? Or it can be any random number?

OpenStudy (abhisar):

yes it always have to be 0, C can be any random number

OpenStudy (yanasidlinskiy):

What would happen if It was a random number?

OpenStudy (abhisar):

like ax^2+bx+c=b ?

OpenStudy (yanasidlinskiy):

Yea. Would It be like an awkward answer?

OpenStudy (abhisar):

ax^+bx+c-b'=0

OpenStudy (abhisar):

it's still 0 on RHS...getting ?

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar like ax^2+bx+c=b' ? \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

Case1 , the left hand side is a perfect square, take root both sides to get the answer Case2, the left hand side is not a perfect square, do as Abhisar did

OpenStudy (yanasidlinskiy):

Wait. how?

OpenStudy (abhisar):

ax^2+bx+c=d (suppose) Then ax^2+bx+c-d=0

OpenStudy (abhisar):

Just bring the right hand term to left

OpenStudy (yanasidlinskiy):

Ooohhh...Ok. I got it:) Thanx, Abhisar!!!!!!:) \(\huge\cal\color{cyan}{Thank~you!!!!}\)

OpenStudy (abhisar):

\(\rlap{\color{red}{\huge\bigstar}\huge \color{red}{ \text{You're Most Welcome! }}\color{red}\bigstar}{\; \color{black}{\huge\bigstar}\huge \color{black}{\text{You're Most Welcome! }}\color{black}\bigstar}\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)